# Accelerating RF Power Amplifier Design via Intelligent Sampling and ML-Based Parameter Tuning

### v1

@@ Abstract
This paper presents  [^1] Submitted to the IEEE International Conference on Future Machine Learning and Data Science (FMLDS 2025). a machine learning-accelerated optimization framework for RF power amplifier design that reduces simulation requirements by 65% while maintaining ±$\pm 0.3-0.4$ dBm accuracy. The proposed method combines MaxMin Latin Hypercube Sampling with CatBoost gradient boosting to intelligently explore multidimensional parameter spaces. Instead of exhaustively simulating all parameter combinations to achieve target P2dB compression specifications, our approach strategically selects approximately 35% of critical simulation points. The framework processes ADS netlists, executes harmonic balance simulations on the reduced dataset, and trains a CatBoost model to predict P2dB performance across the entire design space. Validation across 15 PA operating modes yields an average $R^{2}$ of 0.901, with the system ranking parameter combinations by their likelihood of meeting target specifications. The integrated solution delivers 58.24-77.78% reduction in simulation time through automated GUI-based workflows, enabling rapid design iterations without compromising accuracy standards required for production RF circuits.
@@ III-D CatBoost Model Training and Prediction
The trained model predicts P2dB values for all the remaining parameter combinations, effectively interpolating between the sampled points with ±$\pm 0.3-0.4$ dBm accuracy.
@@ V Results and Validation
The framework achieved remarkable efficiency gains: 65% reduction in required simulations (615 vs. 1755) translated to 58.24-77.78% decrease in total simulation time. This enables engineers to focus on high-probability parameter combinations, achieving 2× speedup in design cycles while maintaining the ±$\pm 0.3-0.4$ dBm accuracy essential for production-standard PA development.

### v2

@@ Abstract
This paper presents  [^1] Submitted to the IEEE International Conference on Future Machine Learning and Data Science (FMLDS 2025). a machine learning-accelerated optimization framework for RF power amplifier design that reduces simulation requirements by 65% while maintaining ±$\pm 0.4$ dBm accuracy for the majority of the modes. The proposed method combines MaxMin Latin Hypercube Sampling with CatBoost gradient boosting to intelligently explore multidimensional parameter spaces. Instead of exhaustively simulating all parameter combinations to achieve target P2dB compression specifications, our approach strategically selects approximately 35% of critical simulation points. The framework processes ADS netlists, executes harmonic balance simulations on the reduced dataset, and trains a CatBoost model to predict P2dB performance across the entire design space. Validation across 15 PA operating modes yields an average $R^{2}$ of 0.901, with the system ranking parameter combinations by their likelihood of meeting target specifications. The integrated solution delivers 58.24-77.78% reduction in simulation time through automated GUI-based workflows, enabling rapid design iterations without compromising accuracy standards required for production RF circuits.
@@ III-D CatBoost Model Training and Prediction
The trained model predicts P2dB values for all the remaining parameter combinations, effectively interpolating between the sampled points with ±$\pm 0.4$ dBm accuracy for the majority of the modes.
@@ V Results and Validation
The framework achieved remarkable efficiency gains: 65% reduction in required simulations (615 vs. 1755) translated to 58.24-77.78% decrease in total simulation time. This enables engineers to focus on high-probability parameter combinations, achieving 2× speedup in design cycles while maintaining the ±$\pm 0.4$ dBm accuracy for the majority of the modes essential for production-standard PA development.